# GeoDANO: Geometric VLM with Domain Agnostic Vision Encoder

### v1

@@ 4.2 Domain adaptation of GeoCLIP
Although GeoCLIP enhances the geometric premises recognition on the benchmark set, the diagram styles in existing PGPS benchmarks differ, necessitating further adaptation. To overcome this challenge, we propose a domain adaptation method for GeoCLIP. To this end, we propose a few-shot domain adaptation method utilizing a few labeled diagrams.
@@
We develop a way to translate the target diagrams into source style. Thankfully, since well-known PGPS datasets come with diagram captions written in formal languages (Lu et al., 2021), we can easily convert them to the AlphaGeometry-style descriptions. Given the translated descriptions, we utilize the rendering engine of AlphaGeometry to translate the target domain images into the source domain. With the translation, we can generate the same diagram in the source domain style. Fig. 6 provides examples of the diagram pairs with different styles. However, in some cases, the original description contains geometric premises that are unrecognizable from the diagram, such as $\angle ACB=35.0$ in Fig. 1(a). Therefore, we apply the same filtering process used in GeoCLIP to translate the AlphaGeometry-style descriptions into natural languages.
@@
We perform domain adaptation by fine-tuning the vision encoder through the style-transferred diagram-caption pairs. Let $\hat{\mathcal{D}}_{T_{j}}$ be a collection of the original diagram and style-transferred captions, i.e., $\hat{\mathcal{D}}_{T_{j}}=\{(D_{T_{j}}^{(i)},\hat{X}_{T_{j}}^{(i)})\}_{i=1}^{N_{T_{j}}}$, and let $\hat{\mathcal{D}}_{T_{j}S}$ be a collection of the original and style transferred diagram pairs, i.e., $\hat{\mathcal{D}}_{T_{j}S}=\{(D_{T_{j}}^{(i)},\hat{D}_{T_{j}}^{(i)})\}_{i=1}^{N_{T_{j}}}$. The cross-domain adaptation objective is written as
@@ 5 Experiments
In this section, we evaluate the PGPS performance of our VLM equipped with the domain adapted GeoCLIP on MathVerse (Zhang et al., 2024a). We compare its performance against established PGPS baselines. We also present ablation studies highlighting our VLM’s strong visual feature recognition and resilience to domain shifts, both of which are facilitated by the adapted vision encoder.
@@ 5.1 Experimental settings and training details
Table 2: PGPS accuracy on MathVerse benchmark. We compare the performance of GeoDANO against PGPS specialist models, which generate a solution program as an output. GeoDANO-OC, -GC, and -GCD are three variants of our model with different encoders. Further details about these variants can be found in Section 5.3.
Datasets.
Training details.
We describe the construction of our geometric VLM with domain-agnostic vision encoder, named GeoDANO. Based on GeoCLIP developed in Section 4.1, we apply the domain adaptation to GeoQA and Geometry3K datasets. For the domain adaptation, we randomly sample 50 diagrams and translate the diagram and caption styles following the procedure described in Section 4.2. Finally, GeoCLIP is fine-tuned via Eq. 2. We name the GeoQA and Geometry3K adapted GeoCLIP as GeoCLIP-DA.
We combine LLama-3-8b-Instruct (Dubey et al., 2024) and GeoCLIP-DA to construct a VLM. The combined model is then fine-tuned again with the training set of GeoQA and PGPS9K to predict the solution program. For PGPS9K, we use the Geometry3K split. While previous works focusing on PGPS do not consider optical character recognition (OCR) from diagrams since the benchmark datasets, GeoQA and PGPS9K, provide necessary details in problem descriptions, numerical values can appear within diagrams in real-world settings. Therefore, we fine-tune GeoDANO with additional OCR capability by modifying the problem statements. Additional details about the modification process with hyper-parameter configurations can be found in Appendix D.
In addition, we unify the programming languages used in the solution programs of GeoQA and PGPS9K by converting GeoQA language into PGPS9K format. The unification makes the output of VLM consistent since both datasets use different types of formal languages.
@@
For the specialist VLMs, we test PGPSNet (Zhang et al., 2023), NGS (Chen et al., 2021), SCA-GPS (Ning et al., 2023), GeoFormer (Chen et al., 2022), UniMath-Flan-T5 (Liang et al., 2023), and GeoX (Xia et al., 2024). For GeoX, we use the two variants GeoX-Geo3K and GeoX-GeoQA, which are fine-tuned on Geometry3K and GeoQA, respectively.
For the generalist VLMs, we test two GPT-4o variants (Hurst et al., 2024): gpt-4o-2024-11-20 and gpt-4o-mini-2024-07-18, and the InternVL2.5 variants: 8B and 26B models (Chen et al., 2024).
@@ 5.2 Results
In Table 2, GeoDANO shows the best performance in almost all the problem variants and metrics except the completion accuracy in the text-dominant task. Note that the specialist models cannot solve the vision-dominant problems since these problems do not contain variables representing numerical values, such as a length, in the problem description. When comparing the performance between text and vision-dominant tasks, the top-10 accuracy of GeoDANO on vision-dominant task is higher than the top-10 accuracy of the specialist models on text-dominant task except for GeoX-GeoQA. Given that the two tasks use the same problem set, the result implies that GeoDANO performs better than the specialist models without having the geometric premises in the problem description. In other words, our vision encoder can extract geometric premises accurately from the visual information.
@@
Table 3: Comparison between GeoDANO and generalist VLMs on multiple choice questions. We report the accuracy of GeoDANO and GeoX following the same evaluation protocol suggested in Zhang et al. (2023).
@@
Table 3 reports the performance of generalist VLMs and GeoDANO on multiple choice questions. GeoDANO outperforms proprietary closed models, i.e., GPT-4o variants, and open-sourced models, i.e., the InternVL2.5 variants. Especially, the performance gap between GeoDANO and InternVL2.5-26B reflects the parameter efficiency of our VLM. While GeoDANO shows impressive results among the variants, the performance of GeoX-GeoQA degrades dramatically as the visual information moves from the text to the diagram. Our work is the first to show that the specialist can compete with the generalist in MathVerse.
@@ 5.3 Ablation studies
We show the experimental result in Table 2. GeoDANO-OC and GeoDANO-GC represent the VLM with OpenCLIP and GeoCLIP without domain adaptation, respectively. GeoDANO-GCD represents the GeoCLIP with additional unfiltered domain captions. GeoDANO outperforms other variants on most tasks, except the completion accuracy on the vision-intensive task.
@@
Figure 3: Visualization of OpenCLIP and GeoCLIP-DA embeddings. The orange, green, and blue dots represent PGPS9K, GeoQA, and synthetic diagrams, respectively. In the top row, the three diagrams on the left and right are those with the highest cosine similarities to the center under OpenCLIP and GeoCLIP-DA, respectively.
@@
We examine how effectively GeoCLIP-DA generalizes to new domains with different diagram styles. For this experiment, we compare the embedding similarity between two diagrams representing the same structure in different styles. To create the paired dataset, we use a similar process described in Section 4.2. Specifically, a total of 100 diagrams are sampled from the test sets of GeoQA and PGPS9K, and these samples are rendered in AlphaGeometry style through the diagram description.
For evaluation, we sample 100 diagrams from each of the target domain’s training sets and compare the similarity against the original diagram via cosine similarity. We also compute the similarity between the style transferred diagram and the original diagram. We report two metrics for test diagrams: the mean rank (MR) and the mean average precision (mAP) of the style-transferred diagram.
As reported in Table 4, GeoCLIP-DA produces similar embeddings for structurally equivalent diagrams, regardless of their stylistic differences. Fig. 3 visualizes the diagram embeddings of OpenCLIP and GeoCLIP-DA. As one can observe, the OpenCLIP embeddings are largely separated by the domain of the diagrams, whereas those of GeoCLIP-DA appear to capture and align with underlying visual features more effectively.
@@ 6 Conclusion
In this work, we present a domain-agnostic VLM for PGPS by refining the vision encoder. Although our VLM exhibits strong performance in recognizing visual features, its coverage remains limited to geometric premises. Building on the success of the synthetic data engine and contrastive learning, extending this combination to different kinds of visual features, e.g., sub-structures in molecular graphs (Kamoi et al., 2024), statistics from charts (Masry et al., 2022), and solid geometry, promises further improvements in recognition of VLM. Due to the limitations in the experimental environment, we are unable to test LLMs with more than 30B parameters.
@@
Appendix A Synthetic Data Engine
Figure 4: Example of randomly sampled AlphaGeometry problems. For each row, the first element describes the randomly sampled AlphaGeometry problem and the others are the geometric premises, diagram, and GeoCLIP-style caption that can be obtained from the AlphaGeometry problem. Note that the GeoCLIP-style caption can be obtained by filtering certain geometric properties, e.g., angle measure, perpendicularity, and concyclicity, from the geometric premises.
@@ Appendix A Synthetic Data Engine
Table 5: Geometric premises used in AlphaGeometry. Visual premises denotes the geometric premises which can be directly perceived from the diagram. Non-visual premises requires reasoning to be recognized.
In this section, we provide the details of our synthetic data engine. Based on AlphaGeometry (Trinh et al., 2024), we generate synthetic diagram and caption pairs by randomly sampling a AlphaGeometry program with Algorithm 1.
@@
Examples for randomly sampled AlphaGeometry problems and their corresponding diagrams and lists of geometric premises are described in Fig. 4. The types of geometric premises that appear in our synthetic data engine are listed in Table 5.
Appendix B Details of Benchmark
B.1 Training details
@@
[…]
B.2 Visualization of the vision encoders
We visualize the embeddings of the vision encoders used in Section 3.2 at Fig. 5.
@@ B.2 Visualization of the vision encoders
Figure 5: The embeddings of the vision encoders on the diagrams of TwoLines task. We visualize the embeddings of the vision encoders on the diagrams of TwoLines task. The blue, orange, and green dots are the diagrams where the two lines AB and BC are collinear, perpendicular, and otherwise, respectively.
Appendix C GeoCLIP-DA
C.1 Domain adaptation data
We adopt GeoCLIP to the two PGPS benchmarks: GeoQA (Chen et al., 2021) and PGPS9K (Zhang et al., 2023). For PGPS9K, we use the Geometry3K split. Fig. 6 shows the pairs used to adapt the domain of GeoCLIP.

### v2

@@ 4.2 Domain adaptation of GeoCLIP
Although GeoCLIP enhances the geometric premises recognition on the benchmark set, the diagram styles in existing PGPS benchmarks differ, necessitating further adaptation. To overcome this challenge, we propose a few-shot domain adaptation method utilizing a few labeled diagrams.
@@
Figure 3: Illustration of the translation process for domain adaptation. We first translate the given geometric diagram from the target domain into an AlphaGeometry problem. We then generate multiple diagrams sharing the same visual geometric premises in the AlphaGeometry style.
We develop a way to translate the target diagrams into the source style. Thankfully, since well-known PGPS datasets come with diagram captions written in formal languages (Lu et al., 2021), we can easily convert them to the AlphaGeometry-style descriptions. Given the translated descriptions, we utilize the rendering engine of AlphaGeometry to translate the target domain images into the source domain. With the translation, we can generate the same diagram in the source domain style. Fig.˜A4 provides examples of the diagram pairs with different styles. However, in some cases, the original description contains geometric premises that are unrecognizable from the diagram, such as $\angle ACB=35.0$ in Fig.˜1(a). Therefore, we apply the same filtering process used in GeoCLIP to translate the AlphaGeometry-style descriptions into natural languages. Additional details in the translation process is described in Section˜D.1 and Fig.˜3.
@@
We adapt the domain by fine-tuning the vision encoder through the style-transferred diagram-caption pairs. Let $\hat{\mathcal{D}}_{T_{j}}$ be a collection of the original diagram and style-transferred captions, i.e., $\hat{\mathcal{D}}_{T_{j}}=\{(D_{T_{j}}^{(i)},\hat{X}_{T_{j}}^{(i)})\}_{i=1}^{N_{T_{j}}}$, and let $\hat{\mathcal{D}}_{T_{j}S}$ be a collection of the original and style transferred diagram pairs, i.e., $\hat{\mathcal{D}}_{T_{j}S}=\{(D_{T_{j}}^{(i)},\hat{D}_{T_{j}}^{(i)})\}_{i=1}^{N_{T_{j}}}$. The cross-domain adaptation objective is written as
@@ 5 Experiments
In this section, we evaluate the PGPS performance of our VLM equipped with the domain-adapted GeoCLIP on MathVerse (Zhang et al., 2024a). We compare its performance against established PGPS baselines. We also present ablation studies highlighting our VLM’s strong visual feature recognition and resilience to domain shifts, both of which are facilitated by the adapted vision encoder.
(a) GeoCLIP-DA
(b) GeoDANO
Figure 4: The overall training process of GeoDANO. The GeoDANO training consists of two stages. (a) Using the CLIP objective, we first train GeoCLIP-DA by optimizing OpenCLIP’s vision and text encoders on synthetic diagram-caption pairs and apply domain adaptation. (b) We then construct GeoDANO by combining GeoCLIP-DA with a projection layer and a language model. We train the projection layer and language model jointly in an end-to-end manner using diagram-text pairs annotated with logic programs.
@@ 5.1 Experimental settings and training details
Datasets.
Training details.
We describe the construction of our geometric VLM with domain-agnostic vision encoder, named GeoDANO. Based on GeoCLIP developed in Section˜4.1, we apply the domain adaptation to GeoQA and Geometry3K datasets. For the domain adaptation, we randomly sample 50 diagrams and translate the diagram and caption styles following the procedure described in Section˜4.2. Finally, GeoCLIP is fine-tuned via Eq.˜2. We name the GeoQA and Geometry3K adapted GeoCLIP as GeoCLIP-DA.
We combine LLama-3-8b-Instruct (Dubey et al., 2024) and GeoCLIP-DA to construct a VLM. The combined model is then fine-tuned again with the training set of GeoQA and PGPS9K to predict the solution program. For PGPS9K, we use the Geometry3K split. Additional details about the training can be found in Appendix˜E.
Modification of training data.
While previous works focusing on PGPS do not consider optical character recognition (OCR) from diagrams since the benchmark datasets, GeoQA and PGPS9K, provide necessary details in problem descriptions, numerical values can appear within diagrams in real-world settings. Specifically, an interesting observation from GeoQA and PGPS9K datasets is that the numerical measurements, such as angles, lengths, and volumes, are not written in the problem description but given as additional conditions, and the numerals are substituted as a variable in the problem description as shown in Fig.˜1(a). Therefore, the VLM only needs to produce the solution program without having optical character recognition (OCR) from the diagram. The variables are automatically substituted with the actual numbers when the program is executed. Therefore, the vision encoders do not need to learn OCR from the image.
However, this approach cannot be generalized to a broader class of problems where the numerals are embedded in the diagram instead of being written in the problem description. Some variants of MathVerse, such as the vision-dominant problems, fall into this category as well. To incorporate OCR into the solution of the problem, we modify some problem statements in the training set, such that the numerical measurements are only shown in the diagram and not in the statements. We further modify the solution problem so that the solution contains OCR results as a part of the final output. Finally, we unify the language of the solution programs used in GeoQA and PGPS9K by converting GeoQA programs into PGPS9K format. The unification makes the output of VLM consistent since both datasets use different types of formal languages. Fig.˜A6 shows examples of the modified input pairs and solutions, where the first problem statement does not have numerical measurements, and the OCR results are in the part of the output solution program.
@@
Table 2: PGPS accuracy on MathVerse benchmark. We compare the performance of GeoDANO against PGPS specialist models, which generate a solution program as an output. GeoDANO-OC, -GC, and -GCD are three variants of our model with different encoders. Further details about these variants can be found in Section˜5.3.
@@
For the specialist VLMs, we test PGPSNet (Zhang et al., 2023), NGS (Chen et al., 2021), SCA-GPS (Ning et al., 2023), GeoFormer (Chen et al., 2022), UniMath-Flan-T5 (Liang et al., 2023), GeoX (Xia et al., 2024), and MAVIS (Zhang et al., 2024b). For GeoX, we use the two variants GeoX-Geo3K and GeoX-GeoQA, which are fine-tuned on Geometry3K and GeoQA, respectively. We mimic MAVIS by replacing the vision encoder of GeoDANO with MAVIS-CLIP, while keeping other components, e.g., the projection layer architecture, language model, and training process, unchanged [^1] Reproducing MAVIS is not feasible due to the unavailability of both the trained model checkpoint and the complete dataset used for training..
For the generalist VLMs, we test two GPT-4o variants (Hurst et al., 2024): gpt-4o-2024-11-20 and gpt-4o-mini-2024-07-18, the InternVL2.5 variants: 8B and 26B models (Chen et al., 2024), SPHINX-MoE (Liu et al., 2024), and Math-PUMA-DeepSeek-Math-VL-7B (Zhuang et al., 2025).
@@ 5.2 Results
In Table˜2, GeoDANO shows the best performance in almost all the problem variants and metrics except the text-dominant task. Note that the specialist models cannot solve the vision-dominant problems since these problems do not contain variables representing numerical values, such as a length, in the problem description. When comparing the performance between text and vision-dominant tasks, the top-10 accuracy of GeoDANO on the vision-dominant task is higher than the top-10 accuracy of the specialist models on the text-dominant task, except for GeoX-GeoQA. Given that the two tasks use the same problem set, the result implies that GeoDANO performs better than the specialist models without having the geometric premises in the problem description. In other words, our vision encoder can extract geometric premises accurately from the visual information.
@@
Table 3: Comparison between GeoDANO and generalist VLMs on multiple choice questions.
@@
Table˜3 reports the performance of generalist VLMs and GeoDANO on multiple choice questions. GeoDANO outperforms proprietary closed models, i.e., GPT-4o variants, and open-sourced models, i.e., the InternVL2.5 variants. Especially, the performance gap between GeoDANO and InternVL2.5-26B reflects the parameter efficiency of our VLM. While GeoDANO shows impressive results among the variants, the performance of GeoX-GeoQA degrades dramatically as the visual information moves from the text to the diagram. Our work is the first to show that the specialist can compete with the generalist in MathVerse.
@@ 5.3 Ablation studies
We show the experimental result in Table˜2. GeoDANO-OC and GeoDANO-GC represent the VLM with OpenCLIP and GeoCLIP without domain adaptation, respectively. GeoDANO-GCD represents the GeoCLIP with additional unfiltered domain captions. GeoDANO outperforms other variants on most tasks, except the completion accuracy on the vision-intensive task.
@@
We examine how effectively GeoCLIP-DA generalizes to new domains with different diagram styles. For this experiment, we compare the embedding similarity between two diagrams representing the same structure in different styles. To create the paired dataset, we use a similar process described in Section˜4.2. Specifically, a total of 100 diagrams are sampled from the test sets of GeoQA and PGPS9K, and these samples are rendered in AlphaGeometry style through the diagram description.
For evaluation, we sample 100 diagrams from each of the target domain’s training sets and compare the similarity against the original diagram via cosine similarity. We also compute the similarity between the style-transferred diagram and the original diagram. We report two metrics for test diagrams: the mean rank (MR) and the mean average precision (mAP) of the style-transferred diagram.
As reported in Table˜4, GeoCLIP-DA produces similar embeddings for structurally equivalent diagrams, regardless of their stylistic differences. Fig.˜A5 visualizes the diagram embeddings of OpenCLIP and GeoCLIP-DA. As one can observe, the OpenCLIP embeddings are largely separated by the domain of the diagrams, whereas those of GeoCLIP-DA appear to capture and align with underlying visual features more effectively.
@@ 6 Conclusion
In this work, we present a domain-agnostic VLM for PGPS by refining the vision encoder. Although our VLM performs strongly in recognizing visual features, its coverage remains limited to geometric premises. Building on the success of the synthetic data engine and contrastive learning, extending this combination to different kinds of visual features, e.g., sub-structures in molecular graphs (Kamoi et al., 2024), statistics from charts (Masry et al., 2022), and solid geometry, promises further improvements in recognition of VLM. Due to the limitations in the experimental environment, we are unable to test LLMs with more than 30B parameters.
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00337955, RS-2023-00217286), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH))
@@ Acknowledgments
Appendix A Details of Related Work
A.1 Comparison between MAVIS and ours
Our contributions differ from MAVIS primarily in three aspects. Firstly, we introduce a systematic benchmark specifically designed for the quantitative analysis of vision encoders’ capabilities in understanding geometric diagrams. This benchmark enables a fine-grained recognition evaluation across distinct geometric features, which is not addressed in MAVIS.
Secondly, we investigate the influence of caption style on vision encoder training by explicitly comparing GeoCLIP-style captions with MAVIS-style captions. MAVIS similarly employs extensive geometric diagram-caption pairs, namely MAVIS-Caption, where the captions include detailed geometric attributes such as object shape and connectivity. However, our empirical results comparing GeoCLIP and GeoCLIP (F ×), where GeoCLIP (F ×) uses captions incorporating all possible geometric premises, demonstrate that redundant geometric information within captions negatively impacts the vision encoder’s recognition performance. Moreover, Table˜1 shows that fine-tuning OpenCLIP with MAVIS-Caption with the setting as MAVIS yields significantly poorer performance on our visual geometric premise recognition benchmark compared to GeoCLIP and OpenCLIP.
Finally, our work explicitly addresses the issue of domain shift across different diagram styles by proposing a few-shot domain adaptation technique, a critical problem not considered in MAVIS.
Appendix B Synthetic Data Engine
Figure A1: Example of randomly sampled AlphaGeometry problems. For each row, the first element describes the randomly sampled AlphaGeometry problem and the others are the geometric premises, diagram, and GeoCLIP-style caption that can be obtained from the AlphaGeometry problem. Note that the GeoCLIP-style caption can be obtained by filtering certain geometric properties, e.g., angle measure, perpendicularity, and concyclicity, from the geometric premises.
@@ Appendix B Synthetic Data Engine
Table A1: Geometric premises used in AlphaGeometry. Visual premises denotes the geometric premises which can be directly perceived from the diagram. Non-visual premises requires reasoning to be recognized.
Figure A2: Illustration of the synthetic diagram-caption pairs generation. We first sample an object from a predefined object set. We then iteratively add points and relations to the existing primitives and relations. Finally, we generate a GeoCLIP-style caption based on the resulting primitives and relations.
In this section, we provide the details of our synthetic data engine. Based on AlphaGeometry (Trinh et al., 2024), we generate synthetic diagram and caption pairs by randomly sampling an AlphaGeometry program with Algorithm˜1. We visualize the AlphaGeometry program and diagram-caption pairs generation process in Fig.˜A2.
@@
Examples for randomly sampled AlphaGeometry problems and their corresponding diagrams and lists of geometric premises are described in Fig.˜A1. The types of geometric premises that appear in our synthetic data engine are listed in Table˜A1.
Appendix C Details of Benchmark
C.1 Role of the textual description
During the evaluation of the vision encoders, only the visual diagram serves as input to the model, and no textual information is provided during training or inference. Specifically, the evaluation is conducted using a linear probing approach, wherein the parameters of the vision encoder remain frozen, and only a linear classifier, initialized randomly, is trained atop the visual embeddings produced by the encoder.
Here, the textual questions from our benchmarks are implicitly represented in the classification labels assigned to each diagram. For instance, the textual question "How are lines AB and BC related?" corresponds directly to classification labels such as "Perpendicular," "Collinear," or "Neither." These labels are used as supervision signals to train the linear classifier. Thus, while the vision encoder receives no explicit textual input, the questions’ semantics are reflected indirectly through the classification labels.
C.2 Training details
@@
We train the linear classifier on the training set of each task for 50 epochs with batch size 128 and learning rate $1\text{e-}4$. We use Adam optimizer for optimization.
C.3 Visualization of the vision encoders
We visualize the embeddings of the vision encoders used in Section˜3.2 at Fig.˜A3.
@@ C.3 Visualization of the vision encoders
Figure A3: The embeddings of the vision encoders on the diagrams of TwoLines task. We visualize the embeddings of the vision encoders on the diagrams of TwoLines task. The blue, orange, and green dots are the diagrams where the two lines AB and BC are collinear, perpendicular, and otherwise, respectively.
Appendix D GeoCLIP-DA
D.1 Details of the translation process
Note that the formal language is mentioned solely as an optional tool to accelerate the translation process when available, but it is not essential. The primary purpose of diagram translation in our method is to generate synthetic diagrams that visually resemble real-world diagrams. Since the objective is visual feature alignment rather than formal semantics, diagrams can effectively be translated manually by simply recreating the visual structure.
While manual translation might appear time-consuming, the effort is minimal and feasible in practice. Specifically, we manually translated only around 50 diagrams per domain, which required less than 3 hours in total. This modest effort substantially improved our model’s cross-domain generalization performance. Thus, manual translation without formal annotations is not only practical but also highly beneficial for domain adaptation. The process of the translation is illustrated in Fig.˜3